# Accepted with Minor Revisions: Value of AI-Assisted Scientific Writing

### v1

@@ 2.4. Peer Review and Scholarly Evaluation
Our experimental design is situated within the broader field of science of peer review, which treats scholarly evaluation not as a purely objective process, but as a human endeavor subject to cognitive and social biases. A significant body of literature has used experimental methods to uncover these biases. For instance, landmark studies have shown that reviewer recommendations are significantly influenced by the prestige of an author’s institution, a bias that double-blinding is intended to mitigate (ceci1982peer; tomkins2017single). Relatedly, we study the perceived origin of the text itself as a bias. The attribution of an abstract to either a human or an AI serves as an analogue to author identity, allowing us to investigate whether a “source identity bias” influences the authorial and review process in a similar manner to established forms of author bias.
In this work, following the methodology of behavioral scientists, we conduct incentivized randomized controlled trials of increasing complexity for this evaluation. Randomized controlled trials are widely regarded as the gold standard for establishing causal relationships in behavioral interventions (banerjee2009experimental). Incentivizing participants in studies ensures higher engagement and more reliable data following a cost-benefit analysis from taking part in the study, thereby yielding results that better reflect real-world usage (gneezy2000pay). Recent studies have employed RCTs to assess the effectiveness of AI tools in various domains, including writing and education (kizilcec2015attrition). These trials provide empirical evidence on how AI interventions influence user outcomes, supporting their utility in scientific research.
Figure 2. Experimental workflow: Authors from a shared pool were randomly assigned to one of four treatments varying abstract source (AI-generated vs. human-written) and information disclosure (with vs. without source information). Authors revised abstracts in the editing interface. Each edited abstract was randomly assigned to three reviewers for their individual verdict. Reviewers see the final edited version of abstracts without knowing what edits were made. A majority vote decides the final verdict.
@@ 3. Experiment Design
Writing scientific prose is a challenging task. It is intended to communicate complex ideas to the academic community or to the public, focusing on preciseness, clarity, and brevity (lindsay2020scientific). It requires the synthesis of diverse specialized knowledge and information sources, including non-textual elements like tables, statistics, and figures, alongside long contexts with strict formatting requirements. Performed by domain experts such as researchers with varying expertise, it is an inherently iterative process (mao2019data; cornell2021please). It is often collaborative, involving people with different experiences contributing, editing, and commenting on each other’s written work.
We aim to create a test-bed to evaluate how humans perceive and evaluate scientific writing. In this work, we focus on the domain of Computer Science. We follow a 4-step approach: (1) We select original research publications from peer-reviewed conference publication venues and curate AI-generated abstracts from the content of these papers using state-of-the-art LLMs; (2) We engineer an incentivized hypothetical conference set-up where we recruit human participants online to form an Author and a Reviewer pool. This allows us to emulate a real-world writing setting for both parties to play their part in the study; (3) We develop a web platform based on FirePad [^5] The editing platform codebase is available at: https://github.com/skai-research/scientific-writing-assistance to capture fine-grained character-level edits such as insertion, deletion, or substitution with timestamps. Each abstract is shown to the authors on this platform to generate edited abstracts. (4) We show each edited abstract to three independent reviewers to obtain the final accept/reject decision by majority voting.
@@ 3.1. Collecting and Curating Abstracts
For the scope of this study, we focus on the computer science (CS) domain for several reasons. First, CS is a fast-paced field with high publication volume (CVPR2024), rapid turnaround cycles, and a high impact factor, particularly in conferences (reuters2024aivstaylor; freyne2010relative). This makes it a relevant and timely setting to study the integration of AI in scientific writing. Second, the widespread adoption of AI tools by researchers in CS provides ecological validity for examining human-AI collaboration in this context. Finally, as CS researchers ourselves, we are well-positioned to assess the accuracy and relevance of domain-specific content. While our study centers on CS, the design is generalizable and can be extended to other scientific domains. A typical research process involves identifying a research problem, reviewing existing literature, formulating a hypothesis, collecting data, analyzing the data, and potentially modifying the hypothesis and repeating the process. At the end, the researchers draw conclusions and present the findings. We evaluate the role of AI-assistance in the final stage, where, given research findings, an LLM generates a coherent abstract summarizing the results of the study in a factual, coherent, and appealing way.
@@
We select 45 published papers from top-tier conference venues in various subfields of CS. [^6] To ensure that our selected research papers were not a part of the training dataset of the state-of-art LLMs, we use only very recent papers published in the top conferences of the domain. For more details, please LABEL:tab:paperlist. For each venue’s latest publication, we randomly sample papers as shown in LABEL:tab:venuelist. Using GPT-4o, we create research excerpts, i.e., extract all relevant text from the paper that is necessary to write an abstract. We provide the exact prompt we use in Figure 13 (Appendix F). Note, we focus on extraction rather than generation to make sure no facts in the original paper are misrepresented in the generated plan. We also manually verify each model output for factual correctness. We provide a sample output in Figure 14 (Appendix F).
Using these excerpts, we also generate versions of AI-generated abstracts with respect to their original counterpart, also using GPT-4o. We chose this model due to its immense popularity. We provide the prompt to generate the abstract and a sample abstract in Figure 15 and Figure 16, respectively (Appendix F). Corresponding original abstracts is provided in Figure 17 (Appendix F).
In this paper, we refer to three versions of abstracts which we use accordingly henceforth. (1) We refer to the authentic human-written abstracts of the published research papers from the peer-reviewed conference publication venues that we used in our study as original abstracts; (2) We refer to the populated AI-generated versions of original abstracts of the respective published papers as AI-generated abstracts; (3) We use the term provided abstracts to indicate those abstracts that we provide the authors to edit. Provided abstracts could be either original abstracts (in Human-noInfo or Human-withInfo treatments) or AI-generated abstracts (in AI-noInfo or AI-withInfo treatments). (4) We use the term edited abstracts to indicate the edited versions of the provided abstracts. To clarify, these edits have been made by the recruited authors. Our investigation includes two participant pools: Authors and Reviewers. We discuss the tasks, incentives, and recruitment procedures of each pool in the respective sections.
@@ 3.2.1. Task
Task 2 entails the main editing task. Every author is presented with three independent research excerpts and their respective provided abstracts. The editing task appears to the authors as illustrated in Figure 9 (Appendix A). Figures 3 and 4 provide the detailed illustration provided to the authors to help them understand the instructions. Authors are instructed to first review the excerpts in the left panel of the screen. A scroll bar is provided to the authors to completely review the excerpt, which may extend beyond the scope of the current screen space. Simultaneously, the right panel entails the editing interface that shows the provided abstract. We explicitly instruct the authors to make all edits in the editing interface only. We disabled copy-pasting from the editing interface so that the participants cannot copy and paste to edit on a platform other than the one provided to them. Additionally, this will also limit authors from generating or making edits using existing AI models. The main goal of the authors is to receive at least two Accept decisions by the reviewers on their edited abstracts. Before starting with Task 2, the authors answer two screening questions (Appendix B). These questions were designed to assess the authors’ understanding of task instructions. Correctly answering these questions was a prerequisite for moving forward with the main editing task.
@@
Authors proceed to Task 3 after they have made their edits and submitted the final versions of edited abstracts. Task 3 involves completing an exit questionnaire that gathers demographic information (such as gender, age, education, current profession), details on how author’s self reported thoughts about the quality and readability of the unedited abstracts, frequency of their edits, kind of edits (such as addition, deletion, correction, substitution or reordering), their familiarity AI-generated text versus human-written text, experience with AI tools in daily life, trust in AI tools, and frequency of AI tool usage for writing tasks.
@@ 3.2.2. Incentives
Authors earn a fixed participation fee of \$15 for participating in this study and completing all the tasks. Note, authors receive this payment irrespective of their abstracts being accepted by the reviewers or not. Using the methodology of economics experiments, we additionally incentivize the participants to receive a bonus pay of \$15 if a randomly selected abstract out of the three edited abstracts receives at least two Accept decisions from the reviewers. Authors also receive a bonus for their answers in the confidence questions, respectively, according to their submitted confidence levels. For example, if an author marks 20 on the slider for the edited abstract and it receives at least two Accept decisions from the reviewers, the authors earn a bonus of \$0.20. An author can receive a maximum of \$0.50 from each of the confidence questions.
To encourage genuine effort and careful editing, our incentive structure followed conventions from experimental economics, where performance-contingent rewards are known to elicit higher attention and task quality (smith1982microeconomic; camerer1999effects). The fixed participation fee ensured baseline engagement, while performance-based bonuses motivated authors to produce edits that improved clarity and persuasiveness rather than completing the task superficially (abeler2011reference). A small confidence-based bonus further encouraged accurate self-assessment, helping us interpret reported confidence as a meaningful behavioral signal (offerman2009truth). Together, these mechanisms help maintain data quality and internal validity while aligning with best practices in incentive-aligned experimental and crowd-sourcing studies (horton2010labor; mason2009financial).
@@ 3.2.3. Procedures and Author Recruitment
We recruit 300 authors via the online crowd-sourcing platform Prolific  [^7] This research has been IRB approved by the Office of Responsible Research Practices at Ohio State University under study number 2024E1034. We also pre-registered for our trial at AEA RCT Registry: https://doi.org/10.1257/rct.16740-1.0 We use the Qualtrics interface to set up our study, which was then linked to Prolific for every author. Participation in the study was voluntary. At the onset of the experiment, each author was shown information about the study, followed by an informed consent form.
For our analysis, we exclude 3 authors owing to malformed responses, resulting in 297 authors. Limiting our study to only the CS domain limits us to restricting authors based on their total number of study approvals on Prolific. The median total approval number for our authors is 336.0 approvals, with a minimum of 0 and a maximum of 5560 approvals. 88.51% of authors have their registered first language as English. The average age bracket for the authors’ pool falls in the range of 25-40, and 42.42% of the authors are female (as self-reported).
Since all our source material for creating the abstracts is strictly limited to the CS domain, we restrict authors to belonging to the CS domain by prescreening them based on their reported Subjects, as recorded on Prolific. It is a general norm in the highly competitive field of CS that students are engaged in scientific writing since their undergraduate years. To optimize for authors to have some experience with formal writing in CS, we also filter authors by their Highest education level completed. Authors include individuals who have an undergraduate degree (BA/BSc/other), a graduate degree (MA/MSc/MPhil/other), and/or a doctorate (PhD/other). 80.00% of authors are currently maintaining their student status, with 55.56% of authors having at least an undergraduate degree with a major/minor in CS.
@@ 3.3. Editing Interface
To build our own interface, we used Firepad, an open-source text editing tool that stores detailed edit history on Firebase Realtime Database and uses CodeMirror as the underlying text editor. While Firepad is no longer being hosted globally, we used its open-source code [^8] https://github.com/FirebaseExtended/firepad to host it ourselves through Firebase Web Hosting. Throughout this study, we used Firepad to record the time, the index, and the edited content per keystroke. We implemented our interface using Firepad 1.5.11, Firebase 5.5.4, and CodeMirror 5.17.0. We plan to open-source this interface for ease of reproduction and use by other researchers. For each author and abstract pair, we create unique Firepad pages. The combination and the index of each pair are used as URL parameters to dynamically populate the authors. Each Firepad page is structured as research findings or excerpts on the left and provided abstracts on the right (Appendix A). To allow authors to refer to the provided abstract when necessary, even though they made changes to the provided abstract, we include the provided abstract at the bottom of the excerpts section. Excerpts are in HTML and embedded in the page and made a non-editable part. Abstracts are located in the editing box. We embed the Firepad links into a Qualtrics [^9] https://www.qualtrics.com/ survey form.
@@ 3.4.1. Task
Reviewers complete three tasks. Task 1 is identical to the Task 1 completed by the authors. Task 2 entails reviewing abstracts. Each reviewer is presented with a set of 20 abstract pairs—the original abstract that accompanied a selected research paper and the respective edited abstract by an author. We inform the reviewers that the edited abstracts encapsulate the same research idea, with the same research methodology and results as the original abstracts. However, it may or may not be expositionally different from the original abstract. Reviewers need to review the edited abstract in reference to the original abstract and answer the following question: Does the edited abstract provide adequate justice to the research idea presented in the original abstract?. The reviewers answer using a slider that ranges from 0 (no justice, worse than the original abstract) to 100 (better than the original abstract), with the default set at 50 (i.e., both abstracts can be considered equally good and one is not better than the other). We also elicit the absolute confidence of the reviewer in each of their 20 decisions. Additionally, to get more insights into the decision-making process of the reviewers, we provide a text box where the reviewers can address this in free-form text. Reviewers proceed to Task 3 after completing Task 2. Task 3 is identical to task 3 as completed by the authors. We ask two additional questions about the publication record of the reviewers and their prior knowledge of the topics discussed in the abstracts.

### v2

@@ 2.4. Peer Review and Scholarly Evaluation
Our experimental design is situated within the broader field of science of peer review, which treats scholarly evaluation not as a purely objective process, but as a human endeavor subject to cognitive and social biases. A significant body of literature has used experimental methods to uncover these biases. For instance, landmark studies have shown that reviewer recommendations are significantly influenced by the prestige of an author’s institution, a bias that double-blinding is intended to mitigate (Ceci and Peters, 1982; Tomkins et al., 2017). Relatedly, we study the perceived origin of the text itself as a bias. The attribution of an abstract to either a human or an AI serves as an analogue to author identity, allowing us to investigate whether a “source identity bias” influences the authorial and review process in a similar manner to established forms of author bias.
In this work, following the methodology of behavioral scientists, we conduct incentivized randomized controlled trials of increasing complexity for this evaluation. Randomized controlled trials are widely regarded as the gold standard for establishing causal relationships in behavioral interventions (Banerjee and Duflo, 2009). Monetary incentives promote task engagement and data reliability. Participants align the study’s rewards with their decisions through a cost-benefit analysis, yielding behavior that closely mirrors real-world decision-making (Gneezy and Rustichini, 2000). Recent studies have employed RCTs to assess the effectiveness of AI tools in various domains, including writing and education (Kizilcec and Halawa, 2015). These trials provide empirical evidence on how AI interventions influence user outcomes, supporting their utility in scientific research.
Figure 2. Experimental workflow: Authors from a shared pool were randomly assigned to one of four treatments varying abstract source (AI-generated vs. human-written) and information disclosure (with vs. without source information). Authors revised abstracts in the editing interface. Each edited abstract was randomly assigned to three reviewers for their individual verdict. Reviewers only see the final edited version of the abstracts. A majority vote decides the final verdict.
@@ 3. Experiment Design
Writing scientific prose is a challenging task. It is intended to communicate complex ideas to the academic community or to the public, focusing on preciseness, clarity, and brevity (Lindsay, 2020). It requires the synthesis of diverse specialized knowledge and information sources, including non-textual elements like tables, statistics, and figures, alongside long contexts with strict formatting requirements. Performed by domain experts such as researchers with varying expertise, it is an inherently iterative process (Mao et al., 2019; Cornell, 2021). It is often collaborative, involving people with different experiences contributing, editing, and commenting on each other’s written work.
We aim to create a test-bed to evaluate how humans perceive and evaluate scientific writing. We follow a 4-step approach: (1) We select original research publications from peer-reviewed conference publication venues and curate AI-generated abstracts from the content of these papers using state-of-the-art LLMs; (2) We engineer an incentivized hypothetical conference set-up where we recruit human participants online to form an Author and a Reviewer pool. This allows us to emulate a real-world writing setting for both parties to play their part in the study; (3) We develop a web platform based on FirePad [^5] The editing platform codebase is available at: https://github.com/skai-research/scientific-writing-assistance to capture fine-grained character-level edits such as insertion, deletion, or substitution with timestamps. Each abstract is shown to the authors on this platform to generate edited abstracts. (4) We show each edited abstract to three independent reviewers to obtain the final accept/reject decision by majority voting.
@@ 3.1. Collecting and Curating Abstracts
For the scope of this study, we focus on the computer science (CS) domain for several reasons. First, CS is a fast-paced field with high publication volume (CVPR 2024, 2024), rapid turnaround cycles, and a high impact factor, particularly in conferences (Reuters, 2024; Freyne et al., 2010). This makes it a relevant and timely setting to study the integration of AI in scientific writing. Second, the widespread adoption of AI tools by researchers in CS provides ecological validity for examining human-AI collaboration in this context. Finally, as CS researchers ourselves, we are well-positioned to assess the accuracy and relevance of domain-specific content. While our study centers on CS, the design is generalizable and can be extended to other scientific domains. A typical research process involves identifying a research problem, reviewing existing literature, formulating a hypothesis, collecting data, analyzing the data, and potentially modifying the hypothesis and repeating the process. At the end, the researchers draw conclusions and present the findings. We evaluate the role of AI-assistance in the final stage, where, given research findings, an LLM generates a coherent abstract summarizing the results of the study in a factual, coherent, and appealing way.
@@
We select 45 published papers from top-tier conference venues in CS. [^6] To ensure that our selected research papers were not a part of the training dataset of the state-of-the-art LLMs, we use a subset of only very recent papers published in 2024 in the top CS conferences. For details on selected papers, please refer to Table 3. We mainly focus on application-based papers from subfields of CS, such as NLP, HCI, Systems, and others. [^7] We refrain from using lengthy papers with theoretical underpinnings, as their heavy dependence on background knowledge may make it challenging to generate concise research excerpts without losing information. For each venue’s latest publication, we randomly sample papers as shown in LABEL:tab:venuelist. Using GPT-4o, we create research excerpts, i.e., extract all relevant text from the paper that is necessary to write an abstract. [^8] OpenAI’s GPT-4o stood as the most recent model available during the time of the experiment. Our conclusions are conditional on this temporal choice and do not presume continued popularity or dominance of the model or its provider. We provide the exact prompt we use in Figure 13 (Appendix). Note, we focus on extraction rather than generation to make sure no facts in the original paper are misrepresented in the generated plan. We also manually verify each model output for factual correctness. We provide a sample output in Figure 14 (Appendix).
Using these excerpts, we also generate versions of AI-generated abstracts with respect to their original counterpart, also using GPT-4o. [^9] At the time of conducting this experiment, ChatGPT’s GPT-4o was OpenAI’s premier offering, with traffic surpassing 100 million visits upon its release (PCMag, 2024; Duarte, 2025). Since its inception, ChatGPT has been in use for writing abstracts, generating and editing text, and academic proofreading (Howard et al., 2024; Al Sawi and Alaa, 2024; Brameier et al., 2023). We provide the prompt to generate the abstract and a sample abstract in Figure 15 and Figure 16 (Appendix). Corresponding original abstract from paper is provided in Figure 17 (Appendix).
In this paper, we refer to four versions of abstracts which we use accordingly henceforth. (1) We refer to the authentic human-written abstracts of the published research papers from the peer-reviewed conference publication venues that we used in our study as original abstracts; (2) We refer to the populated AI-generated versions of original abstracts of the respective published papers as AI-generated abstracts; (3) We use the term provided abstracts to indicate those abstracts that we provide the authors to edit. Provided abstracts could be either original abstracts (in Human-noInfo or Human-withInfo treatments) or AI-generated abstracts (in AI-noInfo or AI-withInfo treatments). (4) We use the term edited abstracts to indicate the edited versions of the provided abstracts. To clarify, these edits have been made by the recruited authors. Our investigation includes two participant pools: Authors and Reviewers. We discuss the tasks, incentives, and recruitment procedures of each pool in the respective sections.
@@ 3.2.1. Task
Task 2 entails the main editing task. Every author is presented with three independent research excerpts and their respective provided abstracts. The editing task appears to the authors as illustrated in Figure 9. Figures 3 and 4 provide the detailed illustration provided to the authors to help them understand the instructions. Authors are instructed to first review the excerpts in the left panel of the screen. A scroll bar is provided to the authors to completely review the excerpt, which may extend beyond the scope of the current screen space. Simultaneously, the right panel entails the editing interface that shows the provided abstract. We explicitly instruct the authors to make all edits in the editing interface only. We disabled copy-pasting from the editing interface so that the participants cannot copy and paste to edit on a platform other than the one provided to them. Additionally, this will also limit authors from generating or making edits using existing AI models. The main goal of the authors is to receive at least two Accept decisions by the reviewers on their edited abstracts. Before starting with Task 2, the authors answer two screening questions (Figure 10). These questions were designed to assess the authors’ understanding of task instructions. Correctly answering these questions was a prerequisite for moving forward with the main editing task.
@@
Authors proceed to Task 3 after they have made their edits and submitted the final versions of edited abstracts. Task 3 involves completing an exit questionnaire that gathers participants’ demographic information (such as gender, education, current profession, age), their first and second-order beliefs about using AI to write abstracts or scientific reports, familiarity AI-generated text versus human-written text, experience with AI tools in daily life, trust in AI tools, and frequency of AI tool usage for writing tasks.
@@ 3.2.2. Incentives
Authors earn a fixed participation fee of \$15 for participating in this study and completing all the tasks. Note, authors receive this payment irrespective of their abstracts being accepted by the reviewers or not. Using the methodology of economics experiments, we additionally incentivize the participants to receive a bonus pay of \$15 if a randomly selected abstract out of the three edited abstracts receives at least two Accept decisions from the reviewers. Authors also receive a bonus for their answers in the confidence questions, respectively, according to their submitted confidence levels. E.g., if an author marks 20 on the slider for the edited abstract and it receives at least two Accept decisions from the reviewers, the authors earn a bonus of \$0.20. An author can receive a maximum of \$0.50 from each of the confidence questions.
To encourage genuine effort and careful editing, our incentive structure followed conventions from experimental economics, where performance-contingent rewards are known to elicit higher attention and task quality (Smith, 1982; Camerer and Hogarth, 1999). The fixed participation fee ensured baseline engagement, while performance-based bonuses motivated authors to produce edits that improved clarity and persuasiveness rather than completing the task superficially (Abeler et al., 2011). This helps us maintain data quality and internal validity while aligning with best practices in incentive-aligned experimental and crowd-sourcing studies (Horton and Chilton, 2010; Mason and Watts, 2009). We do not design confidence questions to be an incentive-compatible belief elicitation. Instead, our bonus serves as a small, performance-contingent incentive to increase effort. Since the acceptance is probabilistic, overstating confidence without producing a high-quality edit does not increase expected payoffs.
@@ 3.2.3. Procedures and Author Recruitment
We recruit 300 authors via the online crowd-sourcing platform Prolific. [^10] This research has been IRB approved by the Office of Responsible Research Practices at Ohio State University under study number 2024E1034. We also pre-registered for our trial at AEA RCT Registry: https://doi.org/10.1257/rct.16740-1.0 We use the Qualtrics interface to set up our study, which was then linked to Prolific for every author. Participation in the study was voluntary. At the onset of the experiment, each author was shown information about the study, followed by an informed consent form.
For our analysis, we exclude 3 authors owing to malformed responses, resulting in 297 authors. Limiting our study to only the CS domain limits us to restricting authors based on their total number of study approvals on Prolific. The median total approval number for our authors is 336.0 approvals, with a minimum of 0 and a maximum of 5560 approvals. 88.51% of authors have their registered first language as English. On average, the authors’ ages (self-reported) range from 25 to 40, and 42.42% of the authors are female (self-reported).
Since all our source material for creating the abstracts is strictly limited to the CS domain, we restrict authors to belonging to the CS domain by prescreening them based on their reported Subjects (proxy for domain of study), as recorded on Prolific. It is a general norm in the highly competitive field of CS that students are engaged in scientific writing since their undergraduate years. To optimize for authors to have some experience with formal writing in CS, we also filter authors by their Highest education level completed. Authors include individuals who have an undergraduate degree (BA/BSc/other), a graduate degree (MA/MSc/MPhil/other), and/or a doctorate (PhD/other). 80.00% of authors are currently maintaining their student status, with 55.56% of authors having at least an undergraduate degree with a major/minor in CS.
@@ 3.3. Editing Interface
To build our own interface, we used Firepad, an open-source text editing tool that stores detailed edit history on Firebase Realtime Database and uses CodeMirror as the underlying text editor. While Firepad is no longer being hosted globally, we used its open-source code [^11] https://github.com/FirebaseExtended/firepad to host it ourselves through Firebase Web Hosting. Throughout this study, we used Firepad to record the time, the index, and the edited content per keystroke. We implemented our interface using Firepad 1.5.11, Firebase 5.5.4, and CodeMirror 5.17.0. We plan to open-source this interface for ease of reproduction and use by other researchers. For each author and abstract pair, we create unique Firepad pages. The combination and the index of each pair are used as URL parameters to dynamically populate the authors. Each Firepad page is structured as research findings or excerpts on the left and provided abstracts on the right (Figure 9). To allow authors to refer to the provided abstract when necessary, even though they made changes to the provided abstract, we include the provided abstract at the bottom of the excerpts section. Excerpts, in HTML, are non-editable and embedded in the page. Provided abstracts are pre-loaded in the editing box. We embed the Firepad links into a Qualtrics [^12] https://www.qualtrics.com/ survey form.
@@ 3.4.1. Task
Reviewers complete three tasks. Task 1 for reviewers is identical to the Task 1 completed by the authors. We elicit reviewers’ perceptions about their LLM use. Task 2 entails reviewing abstracts. Each reviewer is presented with a set of 20 abstract pairs—the original abstract that accompanied a selected research paper and the respective edited abstract by an author. [^13] Showing the original abstract in comparison to the edited versions, eliminates any potential bias in reviewers’ abstract acceptance that may arise from the subjective content of the research. We inform the reviewers that the edited abstracts encapsulate the same research idea, with the same research methodology and results as the original abstracts. However, it may or may not be expositionally different from the original abstract. Reviewers need to review the edited abstract in reference to the original abstract and answer the following question: Does the edited abstract provide adequate justice to the research idea presented in the original abstract? The reviewers answer using a slider that ranges from 0 (no justice, worse than the original abstract) to 100 (better than the original abstract), with the default set at 50 (i.e., both abstracts can be considered equally good and one is not better than the other). We also elicit the absolute confidence of the reviewer in each of their 20 decisions. Additionally, to get more insights into the decision-making process of the reviewers, we provide a text box where the reviewers can address this in free-form text. Reviewers proceed to Task 3 after completing Task 2. Task 3 for reviewers, identical to Task 3 completed by the authors, comprises completing an exit questionnaire. We ask two additional questions about the publication record of the reviewers and their prior knowledge of the topics discussed in the abstracts. Note, we do not disclose the source of the edited abstract to the reviewers. [^14] In our setup, we aim to isolate the reviewers’ perception of the quality of edits in the edited abstracts compared to the original. We do not want to confound acceptance decisions with reviewers’ inherent beliefs about AI-generated scientific abstracts.